algebraic linear equation using the "legal move" concept please help
3x+3+x=7+2x+1+x thank you for any help

Answer by elima(1433)

3x+3+x=7+2x+1+x
Add like terms;
4x+3=3x+8
4x-3x=8-3
x=5
=)